OpenStudy (atomicreaper):

The possible values of x in the equation x^2-3/4x=-1/4 are _____ ± _____ .

OpenStudy (jango_in_dtown):

x^2-3/4x=-1/4 implies x^2-3/4x+1/4=0 Comparing this equation with ax^2+bx+c=0 we have a=1,b=-3/4 and c=1/4 D=b^2-4*a*c=-7/16 so roots are \[(-b \pm \sqrt{D})/2a\] final answer is \[(3/8)\pm (\sqrt{7}/8)i\]